OpenStudy (anonymous):

Factor the Following Polynomials: x²-81 x²-13x+36 Which polynomial is a special product and why?

OpenStudy (anonymous):

does someone want to walk me through these

OpenStudy (anonymous):

apply the formula \[a ^{2}-b ^{2}=(a-b)(a+b)\]

OpenStudy (anonymous):

???

OpenStudy (anonymous):

\[x ^{2}-(9)^{2}=(x-9)(x+9)\]

OpenStudy (anonymous):

how did you get 9

OpenStudy (anonymous):

9^2=81

OpenStudy (anonymous):

oh okay now what

OpenStudy (anonymous):

\[x ^{2}-13x+36=x ^{2}-9x-4x+36\]

OpenStudy (anonymous):

hope you can make factors further......

OpenStudy (anonymous):

I'm still lost idk what work your doing but its not like how its normally done

OpenStudy (anonymous):

Are you not comfortable with factorization methods?

OpenStudy (anonymous):

\[x ^{2}-13x+36=x ^{2}-9x-4x+36=x(x-9)-4(x-9)=(x-4)(x-9)\]